OpenStudy (anonymous):

Evaluate the following expression using the values given: Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3.

OpenStudy (anonymous):

=-2x3-2y2+z4 =-2(3)3-2(5)2+(-3)4 =-2(27)-2(25)+81 =-54-50+81 =81-104 =-23

OpenStudy (anonymous):

@keepsecure93

OpenStudy (anonymous):

thank you <3

OpenStudy (anonymous):

Could you help me with another one of my math problems ?